OpenStudy (anonymous):

Calculate the average rate of change for the given function, from x = −3 to x = 7

OpenStudy (anonymous):

what is the given function, darling

OpenStudy (anonymous):

hold on let me get it up

OpenStudy (anonymous):

x f(x) −3 50 3 10 7 0

OpenStudy (anonymous):

x has -3 3 7 and the rest goes to F(x)

OpenStudy (anonymous):

one fifth negative one fifth −5 5

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

-5 5

jimthompson5910 (jim_thompson5910):

when x = -3, what is the value of f(x) ?

OpenStudy (anonymous):

50

jimthompson5910 (jim_thompson5910):

when x = 7, what is f(x) ?

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

the change in y values is 50-0 = 50 the change in x values is -3-7 = -10 Divide change in y over change in x: 50/(-10) = -5 so the slope, or average rate of change, is -5